Sarah has a collection of nickles, dimes and quarters worth $15.75. She has 10 more dimes than nickles and twice as many quarters as dimes. How many coins of each does she have.
Found 2 solutions by edjones, wvgirl64:
Answer by edjones(8007)

Let n=number of nickles, q=quarters, d=dimes
n=d-10, q=2d
.05n+.1d+.25q=15.75
.05(d-10)+.1d+.25*2d=15.75
.05d-.5+.1d+.5d=15.75
.65d-.5=15.75
.65d-.5+.5=15.75+.5
.65d=16.25
.65d/.65=16.25/.65
d=25
n=15
q=50
.
Ed

Answer by wvgirl64(4)

n=number of nickles, q=quarters, d=dimes
n=d-10, q=2d
.05n+.1d+.25q=15.75
.05(d-10)+.1d+.25*2d=15.75
.05d-.5+.1d+.5d=15.75
.65d-.5=15.75
.65d-.5+.5=15.75+.5
.65d=16.25
.65d/.65=16.25/.65
d=25
n=15
q=50